OpenStudy (anonymous):

Graph the feasible region to the following system of inequalities y-2x≥1 2x+3y≥6

OpenStudy (amistre64):

with what?

OpenStudy (anonymous):

or how would you go about solving it?

OpenStudy (anonymous):

draw the two lines, the feasible region lies above both of the lines

OpenStudy (amistre64):

graph the lines: y-2x = 1 and 2x +3y = 6

OpenStudy (amistre64):

then determine which point, not on the line; satisfies the equatons

OpenStudy (anonymous):

if the inequalities were less than or equal to would i shade below the lines?

OpenStudy (amistre64):

pick (x=0,y=0) since its not on the line.... if it makes a true statement, sahde that side; if false? shaded the other side

OpenStudy (anonymous):

feasible region can be found by putting (0,0) in the line,

OpenStudy (anonymous):

yes amistre is right

OpenStudy (anonymous):

origin is easiest point to check for the required region

OpenStudy (amistre64):

by true statement i mean, use it in the original inequality :)

OpenStudy (anonymous):

\[y-2x \ge 1\]

OpenStudy (anonymous):

substituting x=0 and y=0 o greater than /equal 1,false

OpenStudy (anonymous):

so the area opposite to origin is shaded

OpenStudy (anonymous):

similar for the second line if the two region coincide, it is the required region, otherwise solution does not exist